A brick is dropped from 100 m up. Find its impact velocity and air time


1
Expert's answer
2022-04-26T14:55:19-0400

"v=\sqrt{2gh}=44.3~\frac ms,"

"t=\sqrt{\frac{2h}g}=4.5~s."